# HiSAC: High-Resolution Sensing with Multiband Communication Signals

### v1

@@ 4.1. Intra-subsystem coherent combination
As a first step, we combine the subbands obtained by each subsystem $i$ coherently (since they experience the same TO and PO) to coarsely estimate the multipath delays and complex amplitudes. To do so, we use the OMP algorithm (Eldar and Kutyniok, 2012), which leverages the sparsity of the CIR in the delay domain resulting from the limited number of propagation paths. OMP operates on a grid of candidate path delays, which can be tuned according to the availability of computational resources. We set up a grid with $Q_{i}$ candidates as $0,\dots,(Q_{i}-1)\delta_{i}$, where $\delta_{\rm i}$ is the grid granularity of subsystem $i$. Then, we construct a partial Fourier matrix, $\mathbf{F}_{i}$, that spans all the subcarriers in subsystem $i$ and the delays in the grid. Element $m,q$ of $\mathbf{F}_{i}$ is $\left(\mathbf{F}_{i}\right)_{m,q}=e^{j2\pi mq\delta_{i}\Delta_{f}}/\sqrt{K_{i}}$. Denote by $\mathbf{A}_{i}$ the matrix whose rows are the vectors of all zeros but the $k$-th component, which equals $1$, with $k\in\{k_{i,1},\dots,K_{i,1}-1,\dots,k_{i,S},\dots,K_{i,S}-1\}$. We use $\mathbf{A}_{i}$ to select the rows of $\mathbf{F}_{i}$ whose indices are in the set of available CFR samples in subsystem $i$. Call $\mathbf{\Gamma}_{i}=\mathbf{A}_{i}\mathbf{F}_{i}$, and define $\mathbf{H}_{i}=[H_{i,1,0},\dots,H_{i,1,K_{i,1}-1},\\ \dots,H_{i,S,0},\dots,H_{i,S,K_{i,S}-1}]^{\mathsf{T}}$, that contains all the CFR measurements in subsystem $i$. We denote by $\mathbf{h}_{i}$ the coarse CIR estimate obtained by fusing the $S_{i}$ subbands. The following model holds, $\mathbf{H}_{i}=\mathbf{\Gamma}_{i}\mathbf{h}_{i}+\mathbf{w}_{i}$, where $\mathbf{w}_{i}\sim\mathcal{CN}(0,\sigma^{2}_{w}\mathbf{I})$ is a complex noise vector. Call $||\cdot||_{0}$ the number of non-zero components of a vector. OMP solves the following problem iteratively,
(3) $\mathbf{h}_{i}=\operatorname*{arg\,min}_{\mathbf{h}}||\mathbf{h}||_{0}\mbox{ s. t. }||\mathbf{H}_{i}-\mathbf{\Gamma}_{i}\mathbf{h}||_{2}^{2}<\epsilon,$
@@
to retrieve the sparsest estimate of the CIR that leads to a bounded Mean Squared Error (MSE) with the CFR measurements. The bound on the MSE is regulated by the positive constant $\epsilon$, estimated from the noise level in the CFR estimates. We stop the execution of OMP once the reconstruction error falls below 5% of the norm of $\mathbf{H}_{i}$.
@@
We call $L_{i}^{\rm OMP}$ the number of non-zero components of $\mathbf{h}_{i}$. OMP yields the set of paths delays and complex amplitudes, which we denote by $\{\alpha_{i,l},\tau_{i,l}\}_{l=1}^{L_{i}^{\rm OMP}}$. These correspond to the values and locations of the non-zero elements of vector $\mathbf{h}_{i}$ in the grid. Note that $\tau_{i,l}$ is an estimate of $\tau_{1,l}+\tau_{o,i}$, i.e., the delays $\tau_{i,l}$ contain the relative TO. Finally, a synthetic CFR for subsystem $i$, is obtained as $\tilde{H}_{i,k}=\sum_{l=1}^{L_{i}^{\rm OMP}}\alpha_{i,l}e^{-j2\pi k\Delta_{f}\tau_{i,l}}$, where $k$ can be extended even outside of the subsystems’ bandwidth. This will be used in the next section to compensate for relative TOs and Frequency Offsets.
@@ 4.2. Relative TO and PO compensation
In the second step, the relative TOs and POs among the different ISAC subsystems are compensated for. This is done by leveraging an anchor propagation path between the TX and the RX of each subsystem since the TO and PO are common to all paths (Zhang et al., 2021a). The anchor path could be either the Line-of-Sight (LOS) path, which is commonly assumed to be available in ISAC (Zhang et al., 2022), or a non-LOS static path seen by all subsystems. In the following, we first cast the TO and PO estimation as an optimization problem. Then, we detail how to exploit the anchor path in different subsystems to initialize the TO accurately. Finally, we solve the optimization and compensate for TO and PO to achieve phase-coherence.
Problem formulation. We compensate for $\tau_{o,i}$ and $\phi_{o,i}$ using the synthesized CFR of subsystem $i$ and that of the reference subsystem. TO and PO can be estimated by solving the following minimization problem
@@
where and $\mathrm{Re}\{\cdot\}$ are the complex conjugate and real part of a complex number, respectively. Eq. (4) is high-dimensional, non-linear, and non-convex, which causes solvers to converge to inaccurate solutions and get stuck in local minima. Moreover, its computational complexity is prohibitive since the number of frequency samples $K$ is huge due to its relation with the subcarrier spacing.
@@
Initialization. To solve the problem, we obtain an accurate initial estimate of the TO from the synthesized CFRs. We select the anchor path delay for subsystem $i$, denoted by $\tau_{i,1}$, among $\{\alpha_{i,l},\tau_{i,l}\}_{l=1}^{L_{i}^{\rm OMP}}$. If this corresponds to the LOS, it is easily identifiable by having strong received power compared to scattered paths (Kuschel et al., 2019) and by having the smallest propagation delay. If the anchor path is a non-LOS static reflection, it can be localized by each subsystem before applying HiSAC. For the reference subsystem, $\tau_{1,1}$ is not affected by relative TO. Hence, we estimate the latter as the difference between the anchor path delay of subsystem $i$ and that of the reference subsystem, $\hat{\tau}^{\prime}_{o,i}=\tau_{1,i}-\tau_{1,1}$. This reasoning is shown in Fig. 4(a). Initializing the TO estimate using using delay differences is an innovation of HiSAC. It allows reducing the complexity of directly solving Eq. (4) by reducing the search space for the TO which, unlike the PO, is unbounded and causes fast oscillations of the cost function.
@@
Refinement. Once the initial estimate of the TO has been obtained, we refine it by solving Eq. (4). The computational complexity is greatly reduced by searching over a small neighborhood of $\hat{\tau}^{\prime}_{o,i}$, while for the PO we search over the interval $[0,2\pi]$. We use a grid search for this optimization, focusing the search space in $[\hat{\tau}^{\prime}_{o,i}-\xi,\hat{\tau}^{\prime}_{o,i}+\xi]$ for the TO, where $\xi=5\delta_{i}$, i.e., 5 times the OMP grid step of the subsystem. We use a grid spacing of $1002\xi/$ 100 for TO and $\pi/100$ for PO. As a result, we obtain estimates of the TO, $\hat{\tau}_{o,i}$, and PO, $\hat{\phi}_{o,i}$. Fig. 4(b) shows an example of the refinement step.
@@
TO and PO compensation. The TO and PO are compensated for in each subband by using their estimates as $\bar{H}_{i,s,\kappa}=e^{-j\hat{\phi}_{o,i}}e^{j2\pi\hat{\tau}_{o,i}\kappa\Delta_{f}}H_{i,s,\kappa}$. This enables the coherent combination of the subbands across the full band of interest.
@@ 4.3. Multiband fusion
Once the available subbands have been made mutually coherent, we use OMP to obtain a combined set of delay estimates. Given the large number of subcarriers, $K$, in the total bandwidth, a direct application of OMP would either have prohibitive computational complexity if we select a small grid spacing, or give inaccurate results if the grid spacing is too large. To solve this issue, we leverage the knowledge of the delays estimated by the single subsystems before the coherent fusion to greatly reduce the search space. This can be thought of as focusing the OMP algorithm around the solutions obtained from the lower-resolution subsystems.
Consider the reference subsystem and denote by $\{\tau_{1,l}\}_{l=1}^{L_{1}^{\rm OMP}}$ its set of delays. HiSAC first obtains a union of intervals around the candidate delays from the reference subsystem, i.e., $\mathcal{R}=\bigcup_{l}\mathcal{R}_{l}$, where $\mathcal{R}_{l}=[\tau_{1,l}-\gamma,\tau_{1,l}+\gamma]$. $\gamma$ is chosen as half the best nominal delay resolution among the subsystems, i.e., $\gamma=1/(2\max_{i}(B_{i}))$. Then, HiSAC discretizes $\mathcal{R}$ to construct a grid of $Q$ candidate delays for OMP with step $\delta$. The elements of the discretized set of delays are denoted by $\nu_{1},\dots,\nu_{Q}$. A partial $K\times Q$ Fourier matrix, $\mathbf{F}$ is constructed with the discretized delays, giving $\left(\mathbf{F}\right)_{k,q}=e^{j2\pi k\nu_{q}\Delta_{f}}/\sqrt{K}$. Denote by $\mathbf{A}$ the selector matrix whose rows are the $\mathbf{e}_{k}$, $k\in\{k_{1,1},\dots,K_{1,1}-1,\dots,k_{C,S},\dots,K_{C,S}-1\}$ the matrix that selects the rows of $\mathbf{F}$ whose indices are in the set of available CFR samples. Call $\mathbf{\Gamma}=\mathbf{A}\mathbf{F}\in\mathbb{C}^{M\times Q}$, and define the CFR vector of dimension $M$
@@
that contains all the available measurements from all the $S$ subbands. We denote by $\mathbf{h}\in\mathbb{C}^{Q}$ the CIR obtained by fusing the $S$ subbands. Using OMP we estimate $\mathbf{h}$ by solving the same problem in Eq. (3), using $\bar{\mathbf{H}}$ as the measurement vector and $\mathbf{\Gamma}$ as the model matrix. As for the single subsystems, OMP is stopped once the reconstruction error with respect to the measurements reaches a $5\%$ threshold, and the corresponding number of non-zero components of $\hat{\mathbf{h}}$ is $L^{\rm OMP}$. The set of path delays and amplitudes obtained from the non-zero components of $\hat{\mathbf{h}}$ is $\{\hat{\alpha}_{l},\hat{\tau}_{l}\}_{l=1}^{L^{\rm OMP}}$. The delays are then mapped to relative distances as $\hat{r}_{l}=c\hat{\tau}_{l}-D$, where $D$ is the distance between the TX to the RX, assumed known. Relative distances can be used to localize a target in both mono-static and bi-static scenarios, as described, e.g., in (Kuschel et al., 2019; Pegoraro et al., 2023, 2024).

### v2

@@ 4.1. Intra-subsystem coherent combination
As a first step, we combine the subbands obtained by each subsystem $i$ coherently (since they experience the same TO and PO) to coarsely estimate the multipath delays and complex amplitudes. To do so, we first observe that the CFR in Eq. (1) is sparse in the delay domain. This stems from the fact that in typical communication channels the number of paths, $L$, is much lower than the number of total subcarriers, $K$. Hence, it is possible to represent the CFR as a combination of much fewer basis signals compared to the number of subcarriers. This fact can be exploited to recover the CFR for subsystem $i$ from the incomplete set of $M_{i}$ CFR measurements, according to the compressed sensing principle (Eldar and Kutyniok, 2012).
Notation and definitions. To formulate the intra-subsystem CFR reconstruction as a compressed sensing problem, we set up a grid with $Q_{i}$ channel path candidates as $0,\dots,(Q_{i}-1)\delta_{i}$, where $\delta_{\rm i}$ is the grid granularity of subsystem $i$. The size of the grid should be selected according to a trade-off between reconstruction performance and computational complexity. Each candidate is a complex sinusoidal signal that represents a possible channel path with its corresponding delay. To compactly represent the set of candidate paths, we construct a partial Fourier matrix, $\mathbf{F}_{i}$, that spans all the subcarriers in subsystem $i$ and the delays in the grid. Element $m,q$ of $\mathbf{F}_{i}$ is $\left(\mathbf{F}_{i}\right)_{m,q}=e^{j2\pi mq\delta_{i}\Delta_{f}}/\sqrt{K_{i}}$. The columns of $\mathbf{F}_{i}$ represent the different complex sinusoids corresponding to the candidate paths. However, in our multiband ISAC system not all the subcarriers are observed by subsystem $i$. To model this aspect, we define $\mathbf{A}_{i}$ as the matrix whose rows are the vectors of all zeros but the $k$-th component, which equals $1$, with $k\in\{k_{i,1},\dots,K_{i,1}-1,\dots,k_{i,S},\dots,K_{i,S}-1\}$. We use $\mathbf{A}_{i}$ to select the rows of $\mathbf{F}_{i}$ whose indices are in the set of available CFR samples in subsystem $i$. We obtain the final model matrix, which only contains the available parts of the complex sinusoids corresponding to the candidate paths, as $\mathbf{\Gamma}_{i}=\mathbf{A}_{i}\mathbf{F}_{i}$, of dimension $M_{i}\times K_{i}$. Then, we collect all the available CFR measurements in subsystem $i$ into vector $\mathbf{H}_{i}=[H_{i,1,0},\dots,H_{i,1,K_{i,1}-1},\dots,H_{i,S,0},\dots,H_{i,S,K_{i,S}-1}]^{\mathsf{T}}$, of dimension $M_{i}$.
@@
Compressed sensing problem formulation. To represent the coarse CIR estimate obtained by fusing the $S_{i}$ subbands in subsystem $i$, we use vector $\mathbf{h}_{i}$, of dimension $K_{i}$. $\mathbf{h}_{i}$ is related to the CFR in the $i$-th subsystem by the following model $\mathbf{H}_{i}=\mathbf{\Gamma}_{i}\mathbf{h}_{i}+\mathbf{w}_{i}$, where $\mathbf{w}_{i}$ is a complex Gaussian noise vector with element-wise variance equal to $\sigma^{2}_{w}$. Since the CFR is sparse in the delay domain, only a small fraction of the components of $\mathbf{h}_{i}$ are non-zero.
Estimating the non-zero components of $\mathbf{h}_{i}$ is the aim of the compressed sensing-based CFR reconstruction. Specifically, this can be done by solving the following optimization problem
(3) $\mathbf{h}_{i}=\operatorname*{arg\,min}_{\mathbf{h}}||\mathbf{h}||_{0}\mbox{ subject to }||\mathbf{H}_{i}-\mathbf{\Gamma}_{i}\mathbf{h}||_{2}^{2}<\epsilon,$
where $||\cdot||_{0}$ is the number of non-zero components of a vector and $\epsilon$ is a pre-defined reconstruction error threshold that depends on the noise level.
OMP-based solution. To solve the problem in Eq. (3), we use the OMP algorithm (Eldar and Kutyniok, 2012), which operates by iteratively adding non-zero components to $\mathbf{h}_{i}$ in a greedy fashion. At each iteration, the candidate in $\mathbf{\Gamma}_{i}$ that best correlates with the residual measurements is selected as valid, and its complex amplitude is estimated via least-squares. The new valid candidate is then subtracted from the measurements and the process is repeated.
@@
OMP finds an approximation to the sparsest estimate of the CIR that leads to a bounded Mean Squared Error (MSE) with the CFR measurements. The bound on the MSE is regulated by the positive constant $\epsilon$, which can be estimated from the CFR estimates data. In our implementation, we stop the execution of OMP once the reconstruction error on the original measurements, $\mathbf{H}_{i}$, falls below 5% of the norm of $\mathbf{H}_{i}$.
@@
We call $L_{i}^{\rm OMP}$ the number of non-zero components of $\mathbf{h}_{i}$ as reconstructed by OMP. Together with the complex coefficients of the valid candidate paths, OMP yields the corresponding path delays. We group the coefficients and delays into the set $\{\alpha_{i,l},\tau_{i,l}\}$, $l=1,\dots,L_{i}^{\rm OMP}$. These correspond to the values and locations of the non-zero elements of vector $\mathbf{h}_{i}$ in the grid of candidates. Note that, since no phase offset compensation has been applied yet, $\tau_{i,l}$ is an estimate of $\tau_{1,l}+\tau_{o,i}$, i.e., the delays $\tau_{i,l}$ contain the relative TO. Finally, a synthetic CFR for subsystem $i$, is obtained as $\tilde{H}_{i,k}=\sum_{l=1}^{L_{i}^{\rm OMP}}\alpha_{i,l}e^{-j2\pi k\Delta_{f}\tau_{i,l}},$ where $k$ can be extended even outside of the subsystems’ bandwidth. This will be used in the next section to compensate for relative TOs and POs.
@@ 4.2. Relative TO and PO compensation
In the second step, the relative TOs and POs among the different ISAC subsystems are compensated for. This is done by leveraging an anchor propagation path between the TX and the RX of each subsystem since the TO and PO are common to all paths (Zhang et al., 2021b). The anchor path could be either the LOS path, which is commonly assumed to be available in ISAC (Zhang et al., 2022), or a non-LOS static path seen by all subsystems. In the following, we first cast the TO and PO estimation as an optimization problem. Then, we detail how to exploit the anchor path in different subsystems to initialize the TO accurately. Finally, we solve the optimization and compensate for TO and PO to achieve phase-coherence.
Problem formulation. We compensate for $\tau_{o,i}$ and $\phi_{o,i}$ using the synthesized CFR of subsystem $i$ and that of the reference subsystem. Intuitively, compensating for TO and PO amounts to performing the following two operations: (i) phase-rotating the complex values of the CFR of subsystem $i$, $\tilde{H}_{i,k}$, by $-\phi_{o,i}$ for all subcarriers, and (ii) re-modulating the CFR by multiplying it by the complex exponential $e^{j2\pi k\Delta_{f}\tau_{o,i}}$. It can be seen that applying (i) and (ii) cancels out the offsets from the CFR in Eq. (2).
As a result, TO and PO can be estimated by solving the following minimization problem
@@
where and $\mathrm{Re}\{\cdot\}$ are the complex conjugate and real part of a complex number, respectively. Solving Eq. (4) gives the phase rotation and re-modulation delay that best match the measured CFR on the reference subsystem. The problem is high-dimensional, non-linear, and non-convex, which causes solvers to converge to inaccurate solutions and get stuck in local minima. Moreover, its computational complexity is prohibitive since the number of frequency samples $K$ is huge due to its relation with the subcarrier spacing.
@@
Initialization. To solve the problem, we obtain an accurate initial estimate of the TO from the synthesized CFRs. We select the anchor path delay for subsystem $i$, denoted by $\tau_{i,1}$, among $\{\alpha_{i,l},\tau_{i,l}\},l=1,\dots,L_{i}^{\rm OMP}$. If this corresponds to the LOS, it is easily identifiable by having strong received power compared to scattered paths (Kuschel et al., 2019) and by having the smallest propagation delay. If the anchor path is a non-LOS static reflection, it can be localized by each subsystem before applying HiSAC. The path delays of the reference subsystem, $\tau_{1,1}$, are not affected by relative TO. Since the TO is common to all the propagation paths, it can be estimated from the anchor path and used to initialize the solution to Eq. (4). We estimate the relative TO as the difference between the anchor path delay of subsystem $i$ and that of the reference subsystem, $\hat{\tau}^{\prime}_{o,i}=\tau_{1,i}-\tau_{1,1}$. This reasoning is shown in Fig. 4(a). The OMP grid step used in subsystem $i$, $\delta_{i}$, limits the accuracy of $\hat{\tau}^{\prime}_{o,i}$, since the values of $\tau_{1,i}$ can only lie on the grid. Therefore, we apply a refinement step to the TO estimate and also obtain the PO.
Initializing the TO estimate using delay differences is an innovation of HiSAC. It allows reducing the complexity of directly solving Eq. (4) by reducing the search space for the TO which, unlike the PO, is unbounded and causes fast oscillations of the cost function. Moreover, note that solving Eq. (4) requires achieving phase coherence between the recostructed CFRs across the full bandwidth. When the frequency bands of the subsystems are widely separated, this is challenging since even a small error on the TO can cause large phase variations across wide frequency bands. Existing methods based on linear fitting of the phase, e.g., (Kotaru et al., 2015), are prone to errors in these cases as we show in Section 6.
@@
Refinement. Once the initial estimate of the TO has been obtained, we refine it by solving Eq. (4). The computational complexity is greatly reduced by searching over a small neighborhood of $\hat{\tau}^{\prime}_{o,i}$, while for the PO we search over the interval $[0,2\pi]$. We use a grid search for this optimization, focusing the search space in $[\hat{\tau}^{\prime}_{o,i}-\xi,\hat{\tau}^{\prime}_{o,i}+\xi]$ for the TO, where $\xi=5\delta_{i}$, i.e., 5 times the OMP grid step of the subsystem. Note that this choice achieves a trade-off between reducing the search space as much as possible and accounting for errors in the initialization, whose accuracy depends on $\delta_{i}$. We use a grid spacing of $1002\xi/$ 100 for TO and $1002\pi/$ 100 for PO, thus searching over a $100\times 100$ grid. The size of the grid can be reduced or increased depending on computational resources. As a result, we obtain estimates of the TO, $\hat{\tau}_{o,i}$, and PO, $\hat{\phi}_{o,i}$. Fig. 4(b) shows an example of the refinement step. The initialization of the TO allows restricting the search to a small area in the cost function that presents a clear minimum.
TO and PO compensation. The TO and PO are compensated for in each subband by applying the phase-rotation and re-modulation steps (see points (i) and (ii) in the problem formulation) using the estimates $\hat{\tau}_{o,i}$, and $\hat{\phi}_{o,i}$. This is mathematically expressed as
(5) $\bar{H}_{i,s,\kappa}=e^{-j\hat{\phi}_{o,i}}e^{j2\pi\hat{\tau}_{o,i}\kappa\Delta_{f}}H_{i,s,\kappa},$
@@
where the resulting CFR for subband $s$ of subsystem $i$, $\bar{H}_{i,s,\kappa}$, is now phase-coherent with the reference subsystem, up to estimation errors. This enables the coherent combination of the subbands across the full band of interest.
@@ 4.3. Multiband fusion
Once the available subbands have been made mutually coherent, we use OMP to obtain a combined set of delay estimates. This step is similar to the intra-subsystem CFR reconstruction in Section 4.1. However, the search space is huge since the number of subcarriers in the full band is much larger than that in the single subsystems, i.e., $K\gg K_{i}$. Hence, a direct application of OMP would either have prohibitive computational complexity if we select a small grid spacing, or give inaccurate results if the grid spacing is too large. To solve this issue, we leverage the knowledge of the delays estimated by the single subsystems before the coherent fusion to greatly reduce the search space. This can be thought of as focusing the OMP algorithm around the solutions obtained from the lower-resolution subsystems.
Search space reduction. To effectively reduce the search space, we consider the reference subsystem and denote by $\{\tau_{1,1},\dots,\tau_{1,L_{1}^{\rm OMP}}\}$ its set of delays. Instead of considering a complete grid of delays, as done in the intra-subsystem CFR reconstruction, here HiSAC works on a union of discrete intervals around the reference subsystem’s solutions. To do so, we first obtain a continuous union of intervals around the candidate delays from the reference subsystem, i.e., $\mathcal{R}=\bigcup_{l}\mathcal{R}_{l}$, where $\mathcal{R}_{l}=[\tau_{1,l}-\gamma,\tau_{1,l}+\gamma]$. $\gamma$ is the one-sided width of the interval and it is chosen as half the worst nominal delay resolution among the subsystems, i.e., $\gamma=1/(2\min_{i}(B_{i}))$. This way, the search area around the subsystem’s solution is conservatively adapted to the subsystem with the lowest resolution.
Then, HiSAC discretizes $\mathcal{R}$ to construct a grid of $Q$ candidate delays for OMP with step $\delta$. The latter determines the final HiSAC delay resolution and accuracy. The elements of the discretized set of delays are denoted by $\nu_{1},\dots,\nu_{Q}$.
Thanks to the search space reduction step, the dimensionality of the problem is reduced from having $K$ candidates to $Q$ candidates. $Q$ can be significantly smaller than $K$, since $L_{1}^{\rm OMP}$ is small due to the sparsity of the channel and $\delta$ is under our control.
OMP-based solution. Similar to Section 4.1, the subsystem combination can be cast as a compressed sensing problem that we solve using OMP. To this end, we construct a matrix with the candidate complex sinusoidal components in the channel as its columns. This is a partial $K\times Q$ Fourier matrix, $\mathbf{F}$, that only contains the candidates corresponding to the delays in set $\{\nu_{1},\dots,\nu_{Q}\}$. The components of $\mathbf{F}$ are defined as $\left(\mathbf{F}\right)_{k,q}=e^{j2\pi k\nu_{q}\Delta_{f}}/\sqrt{K}$. To model the missing subcarrier measurements, we introduce a selector matrix, $\mathbf{A}\in\{0,1\}^{M\times K}$, whose rows are the $K$-dimensional vectors of all zeros but the $k$-th component, and each row has a different $k\in\{k_{1,1},\dots,K_{1,1}-1,\dots,k_{C,S},\dots,K_{C,S}-1\}$. Left-multiplying $\mathbf{A}$ by $\mathbf{F}$ selects the rows of $\mathbf{F}$ whose indices are in the set of available CFR samples. Call $\mathbf{\Gamma}=\mathbf{A}\mathbf{F}\in\mathbb{C}^{M\times Q}$, and define the CFR vector of dimension $M$
@@
that contains all the available measurements from all the $S$ subbands after TO and PO compensation. We denote by $\mathbf{h}\in\mathbb{C}^{Q}$ the CIR obtained by fusing the $S$ subbands. Using OMP we estimate $\mathbf{h}$ by solving the same problem in Eq. (3), using $\bar{\mathbf{H}}$ as the measurement vector and $\mathbf{\Gamma}$ as the model matrix. As for the single subsystems, OMP is stopped once the reconstruction error with respect to the measurements reaches a $5\%$ threshold, and the corresponding number of non-zero components of $\hat{\mathbf{h}}$ is $L^{\rm OMP}$. The set of path delays and amplitudes obtained from the non-zero components of $\hat{\mathbf{h}}$ is $\{\hat{\alpha}_{l},\hat{\tau}_{l}\},l=1,\dots,L^{\rm OMP}$. The delays are then mapped to relative distances as $\hat{r}_{l}=c\hat{\tau}_{l}-D$, where $D$ is the distance between the TX and the RX, assumed known. Relative distances can be used to localize a target in both mono-static and bi-static scenarios, as described, e.g., in (Kuschel et al., 2019; Pegoraro et al., 2024b, a).